The sum of Alex's age and that of his father is 56. Three years before, Alex's age was a third of his father's age. In how many years Alex's age will be one third that of his father's age?
Answer by mananth(16946)

let Alex's age be x
father's age = 56-x
...
3 years ago
their ages were
son = x-3
father = 56-x-3 = 53-x
...
x-3 = 1/3 * (53-x)
multiply by 3
3(x-3)=53-x
3x-9 = 53-x
3x+x=53+9
4x=62
/4
x= 15.5 years son's age
father's age = 40.5 years
..
m.ananth@hotmail.ca